OpenStudy (anonymous):

When 20.0 g C2H6 and 60.0 g O2 react to form CO2 and H2O, how many grams of water are formed?

OpenStudy (cuanchi):

2C2H6+7O2->4CO2+6H2O \[20g C _{2}H _{6}*\frac{6*MM _{H _{2}O}}{ 2*MM _{C _{2}H _{6}} }= x_{1} gH _{2} O\] \[60g O _{2}*\frac{6*MM _{H _{2}O}}{ 7*MM _{O_{2}} }= x_{2} gH _{2} O\] MM H2O = 18g MM C2H6 = 30.1g Whichever is smaller x1 or x2 that is the amount of H2O. I guess the O2 it is in excess so it is going to be x1=35g of H2O